# Embedding Space Allocation with Angle-Norm Joint Classifiers for Few-Shot Class-Incremental Learning

### v1

@@ Abstract
Few-shot class-incremental learning (FSCIL) aims to continually learn new classes from only a few samples without forgetting previous ones, requiring intelligent agents to adapt to dynamic environments. FSCIL combines the characteristics and challenges of class-incremental learning and few-shot learning: (i) Current classes occupy the entire feature space, which is detrimental to learning new classes. (ii) The small number of samples in incremental rounds is insufficient for fully training. In existing mainstream virtual class methods, for addressing the challenge (i), they attempt to use virtual classes as placeholders. However, new classes may not necessarily align with the virtual classes. For the challenge (ii), they replace trainable fully connected layers with Nearest Class Mean (NCM) classifiers based on cosine similarity, but NCM classifiers do not account for sample imbalance issues. To address these issues in previous methods, we propose the class-center guided embedding Space Allocation with Angle-Norm joint classifiers (SAAN) learning framework, which provides balanced space for all classes and leverages norm differences caused by sample imbalance to enhance classification criteria. Specifically, for challenge (i), SAAN divides the feature space into multiple subspaces and allocates a dedicated subspace for each session by guiding samples with the pre-set category centers. For challenge (ii), SAAN establishes a norm distribution for each class and generates angle-norm joint logits. Experiments demonstrate that SAAN can achieve state-of-the-art performance and it can be directly embedded into other SOTA methods as a plug-in, further enhancing their performance.
@@
few-shot learning , incremental learning , embedding space allocation , nearest class mean
@@
[label1]organization=State Key Laboratory for Novel Software Technology, country=China
@@ 1 Introduction
Figure 1: The motivation of SAAN. In only CE, old categories fill the entire feature space, making it hard for new categories to be inserted without overlapping. In SAAN, the old categories are allocated to a limited space, guided by the category centers, allowing new categories to be inserted into the reserved space without overlapping. In addition, compared to NCM, which has only angular decision boundaries, SAAN has both angular and norm boundaries, which fully utilizes the feature information to enhance the basis for discrimination.
The goal of FSCIL is to acquire new knowledge effectively while preventing the forgetting of previously learned knowledge. Previous research on FSCIL shows that models trained incrementally are prone to overfit on limited new data and forget the old knowledge catastrophically [14, 15, 8]. To combat overfitting and catastrophic forgetting, some studies propose the use of a feature extractor freezing paradigm [9, 10, 11, 16]. This paradigm freezes most layers in the Convolutional Neural Network (CNN), which is the feature extractor, and uses the Nearest Class Mean (NCM) algorithm [17] instead of Fully Connected (FC) layers as the classifier during the incremental phase. According to the neural collapse phenomenon [18, 19], researchers discover that without any intervention, current training samples would occupy the entire embedding space. This phenomenon is beneficial for enhancing the neural network’s capabilities in conventional classification tasks. However, in the context of Class-Incremental Learning (CIL), learning new classes becomes extremely challenging. This phenomenon addresses the challenge (i) we aim to solve. To address this issue, some researchers [16, 11] propose occupying some of the embedding space with virtual classes generated through data augmentation. Then, during the incremental phase, the space occupied by these virtual classes can be used for inserting new classes. However, these methods cannot control the location of the space occupied by virtual classes, as the occupied positions are determined by the features of the virtual samples. Some researchers [12] propose using pre-allocated fixed prototypes to reduce the feature misalignment problem during the incremental process. However, this approach of completely fixed and randomly allocated prototypes ignores the intrinsic meaning of the samples, which may lead to a discrepancy between the semantic distance of the samples and their distance in the feature space.
@@
CCSA is designed to address the challenge (i) posed by CIL. Specifically, CCSA divides the feature space into multiple subspaces and allocates different subspaces for different incremental sessions by setting class centers, using our designed Cosine Center (CC) loss. CCSA resolves two issues found in previous work. First, in prior virtual class methods [11, 16], due to the lack of guidance for new classes, the reserved space for virtual classes may not necessarily be utilized by new classes. CCSA imposes explicit spatial constraints on new classes, ensuring that they enter the reserved space. Second, unlike methods with completely fixed prototype allocation [12, 23], where random allocation may cause similar classes to be far apart in the feature space, SAAN enables the adjustment of class centers within the designated space to ensure that prototype similarity aligns with the samples’ semantic information.
@@
Since SAAN is designed to address issues in previous virtual class methods, it can be directly integrated into these State-Of-The-Art (SOTA) methods as a plug-in. Experiments show that SAAN can achieve performance close to SOTA and can further improve the performance of existing SOTA methods.
@@
1) CCSA guides sample points into specific subspaces through our meticulously designed CC loss. This resolves two issues found in previous methods: new classes may fail to appear in the reserved feature space, and similar classes may be far apart in the feature space.
@@ 2 Related Works
Few-Shot Class-Incremental Learning. FSCIL diverges from CIL in its requirement for learning from limited samples for novel classes. FSCIL is primarily challenged by the catastrophic forgetting of previously learned classes and a propensity for overfitting to new classes. To enable the model to have forward-compatible learning capabilities, some researchers [11, 16] create virtual classes to reserve space for future classes, while others [12, 23] set fixed prototypes for all classes in advance. FACT [11, 23] proposes a forward compatible training by assigning virtual prototypes to squeeze the embedding of known classes and reserving for new classes. Besides, FACT forecasts potential new categories and prepares for the updating process. The virtual prototypes enable the model to accommodate future updates, serving as proxies dispersed throughout the embedding space to enhance the classifier’s performance during inference. SAVC [16] innovates by distinguishing new classes from existing ones through the incorporation of virtual classes into supervised contrastive learning (SCL). The virtual classes, generated through predefined transformations, serve not only as placeholders for unseen categories within the representation space but also enrich it with diverse semantic information. These virtual class methods demonstrate the importance of reserving embedding space for new classes and enhancing the separability of embeddings.
NC-FSCIL [12] sets up classifiers with Equiangular Tight Frame (ETF) structure that considers future classes, and it induces class alignment with the corresponding classifier to mitigate the classifier misalignment issue during the incremental process. OrCo [23] maximizes the boundaries between classes and reserves space for subsequent incremental data by introducing perturbed prototypes in the feature space and maintaining orthogonality between classes. These methods of pre-allocated prototypes highlight the importance of prior space allocation and preventing misalignment of old categories in new sessions.
@@ 3.1 Problem Setting
In FSCIL, the model is fed with a sequence of training dataset $\left\{\mathcal{D}_{train}^{t}\right\}_{t=0}^{T}$, where $\mathcal{D}_{train}^{t}$ is the training set at the session $t$ and $T$ is the number of all sessions. $\mathcal{D}_{train}^{t}=\left\{(\mathbf{x}_{i},y_{i})\right\}$, where $\mathbf{x}_{i}$ is the $i$-th training sample and $y_{i}$ is the $i$-th label. $\left|\mathcal{D}_{train}^{t}\right|$ denotes the number of the training samples at session $t$. Each session $t$ has a distinct label space $\mathcal{Y}^{t}$, which means that $\mathcal{Y}^{t}\cap\mathcal{Y}^{t^{\prime}}=\varnothing$ for any $t\neq t^{\prime}$. $\left|\mathcal{Y}^{t}\right|$ denotes the number of the classes at session $t$. In the base session, $\left|\mathcal{D}_{train}^{0}\right|$ is equal to the number of all the samples whose labels $\in\mathcal{Y}^{0}$ in the dataset. For example, on CIFAR100 [48], there are 500 training samples for each class at the base session. But, at the incremental sessions, we only can access a limited number of training samples. The training data is organized in an $N$-way $K$-shot format which means there are $N$ classes, and each class contains $K$ training samples at each incremental session. For example, on CIFAR100, each incremental session includes 5 classes, with each class having 5 samples, meaning $N=5$ and $K=5$. At each session $t$, the model is tested on all test samples whose labels $\in(\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t})$.
@@ 3.2 The Model Framework
where ${n^{t}_{j}}$ is the number of samples of class $j$ and $\mathbb{I}\left(\cdot\right)$ is the indicator function. As indicated by the Eq. 1, the NCM classifier, which classifies based on the nearest neighbor principle by measuring the cosine similarity between sample embeddings and representative points, completely disregards the norm information of the sample embeddings.
@@ 3.3 Learning of the Baseline Model
The optimization goal of the feature extractor freezing paradigm baseline is to minimize $\mathcal{L}_{ce}(\phi(\mathbf{x}),y)$, where $\phi(\mathbf{x})=W^{\top}f(\mathbf{x})$ is the model composed of the feature extractor and the linear classifier, and $\mathcal{L}_{ce}(\cdot)$ is the CE loss function. Here, $\phi(\mathbf{x})\in\mathbb{R}^{|\mathcal{Y}|\times 1}$, $W\in\mathbb{R}^{d\times|\mathcal{Y}|}$, and $f(\mathbf{x})\in\mathbb{R}^{d\times 1}$. After the base session, most layers of the CNN are frozen and fine-tuned. The NCM classifier expands by $W=\{\boldsymbol{w}^{0}_{1},\boldsymbol{w}^{0}_{2},\cdots,\boldsymbol{w}^{t}_{|\mathcal{Y}^{t}|-1},\boldsymbol{w}^{t}_{|\mathcal{Y}^{t}|}\}$. $\boldsymbol{w}^{t}$ is continuously expanded during the incremental process based on new data. Although the feature-freezing base approach is simple, it outperforms many FSCIL methods, making it a strong baseline model.
Figure 2: Illustration of CCSA. CCSA sets orthogonal class centers, and assigns these centers to different categories to allocate the embedding space. CCSA uses the well-designed CC loss to guide embeddings into the corresponding subspace, achieving the effect of feature alignment. The centers are adjusted based on semantics in the base session, and new inserted classes are allocated to the remaining space during the incremental sessions.
@@ 4.1 Class-Center Guided Embedding Space Allocation
The baseline model trained solely with CE loss focuses only on the current sample during training and utilizes the entire embedding space, rendering it challenging for the model to assimilate new class knowledge [18, 19]. Inspired by the center loss [49], we mitigate this challenge by establishing learnable class centers (prototypes) and guiding sample embeddings towards their corresponding class centers. Unlike center loss, CCSA constrains the allocation of center positions and the update of centers to achieve the goal of constraining the model’s utilization of the embedding space and providing balanced space for new classes. The overall process of CCSA is shown in Fig. 2.
@@ 4.1.1 Space allocation
We first set $(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)$ orthogonal class centers and assign them one-to-one based on their distance from the class means. To ensure that the total distance between all class centers and their corresponding class means is minimized, we use the Hungarian algorithm [50, 51] to assign the class centers to different categories. The Hungarian algorithm is used to solve the assignment problem in a bipartite graph, aiming to find an optimal matching that minimizes the total cost. We define an $(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)\times(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)$ cost matrix $D=[d_{i,j}]$, where $d_{i,j}$ represents the cosine distance from $i$-th class center to the $j$-th class. Our objective is to find a one-to-one matching that minimizes the total cost. Mathematically, this can be written as:
$\min_{\sigma}\sum_{i=1}^{n}d_{i,\sigma(i)},$

### v2

@@ Abstract
Few-shot class-incremental learning (FSCIL) aims to continually learn new classes from only a few samples without forgetting previous ones, requiring intelligent agents to adapt to dynamic environments. FSCIL combines the characteristics and challenges of class-incremental learning and few-shot learning: (i) Current classes occupy the entire feature space, which is detrimental to learning new classes. (ii) The small number of samples in incremental rounds is insufficient for fully training. In existing mainstream virtual class methods, to address the challenge (i), they attempt to use virtual classes as placeholders. However, new classes may not necessarily align with the virtual classes. For challenge (ii), they replace trainable fully connected layers with Nearest Class Mean (NCM) classifiers based on cosine similarity, but NCM classifiers do not account for sample imbalance issues. To address these issues in previous methods, we propose the class-center guided embedding Space Allocation with Angle-Norm joint classifiers (SAAN) learning framework, which provides balanced space for all classes and leverages norm differences caused by sample imbalance to enhance classification criteria. Specifically, for challenge (i), SAAN divides the feature space into multiple subspaces and allocates a dedicated subspace for each session by guiding samples with the pre-set category centers. For challenge (ii), SAAN establishes a norm distribution for each class and generates angle-norm joint logits. Experiments demonstrate that SAAN can achieve state-of-the-art performance and it can be directly embedded into other SOTA methods as a plug-in, further enhancing their performance.
@@
few-shot learning , incremental learning , embedding space allocation , prototype learning
@@
[label1]organization=National Key Laboratory for Novel Software Technology, country=China
@@ 1 Introduction
Figure 1: The motivation of SAAN. In only cross-entropy supervision, old categories fill the entire feature space, making it hard for new categories to be inserted without overlapping. In SAAN, the old categories are allocated to a limited space, guided by the category centers, allowing new categories to be inserted into the reserved space without overlapping. In addition, compared to NCM, which has only angular decision boundaries, SAAN has both angular and norm boundaries, which fully utilizes the feature information to enhance the basis for discrimination.
The goal of FSCIL is to acquire new knowledge effectively while preventing the forgetting of previously learned knowledge. Previous research on FSCIL shows that models trained incrementally are prone to overfit on limited new data and forget the old knowledge catastrophically [14, 15, 8]. To combat overfitting and catastrophic forgetting, some studies propose the use of a feature extractor freezing paradigm [9, 10, 11, 16]. This paradigm freezes most layers in the Convolutional Neural Network (CNN), which is the feature extractor, and uses the Nearest Class Mean (NCM) algorithm [17] instead of fully connected layers as the classifier during the incremental phase. According to the neural collapse phenomenon [18, 19], researchers discover that without any intervention, current training samples would occupy the entire embedding space. This phenomenon is beneficial for enhancing the neural network’s capabilities in conventional classification tasks. However, in the context of Class-Incremental Learning (CIL), learning new classes becomes extremely challenging. This phenomenon addresses the challenge (i) we aim to solve. To address this issue, some researchers [16, 11] propose occupying some of the embedding space with virtual classes generated through data augmentation [20, 21]. Then, during the incremental phase, the space occupied by these virtual classes can be used for inserting new classes. However, these methods cannot control the location of the space occupied by virtual classes, as the occupied positions are determined by the features of the virtual samples. Some researchers [12] propose using pre-allocated fixed prototypes to reduce the feature misalignment problem during the incremental process. However, this approach of completely fixed and randomly allocated prototypes ignores the intrinsic meaning of the samples, which may lead to a discrepancy between the semantic distance of the samples and their distance in the feature space.
@@
CCSA is designed to address the challenge (i) posed by CIL. Specifically, CCSA divides the feature space into multiple subspaces and allocates different subspaces for different incremental sessions by setting class centers, using our designed cosine center loss. CCSA resolves two issues found in previous work. First, in prior virtual class methods [11, 16], due to the lack of guidance for new classes, the reserved space for virtual classes may not necessarily be utilized by new classes. CCSA imposes explicit spatial constraints on new classes, ensuring that they enter the reserved space. Second, unlike methods with completely fixed prototype allocation [12, 25], where random allocation may cause similar classes to be far apart in the feature space, SAAN enables the adjustment of class centers within the designated space to ensure that prototype similarity aligns with the samples’ semantic information.
@@
Since SAAN is designed to address issues in previous virtual class methods, it can be directly integrated into these State-Of-The-Art (SOTA) methods as a plug-in. Experiments in both conventional and open-ended scenarios demonstrate that SAAN not only achieves performance comparable to SOTA but also further enhances the performance of existing SOTA methods.
@@
1) CCSA guides sample points into specific subspaces through our meticulously designed cosine center loss. This resolves two issues found in previous methods: new classes may fail to appear in the reserved feature space, and similar classes may be far apart in the feature space.
@@ 2 Related Works
Few-Shot Class-Incremental Learning. FSCIL diverges from CIL in its requirement for learning from limited samples for novel classes. FSCIL is primarily challenged by the catastrophic forgetting of previously learned classes and a propensity for overfitting to new classes. To enable the model to have forward-compatible learning capabilities, some researchers [11, 16] create virtual classes to reserve space for future classes, while others [12, 25] set fixed prototypes for all classes in advance. FACT [11, 25] proposes a forward-compatible training by assigning virtual prototypes to squeeze the embedding of known classes and reserving for new classes. Besides, FACT forecasts potential new categories and prepares for the updating process. The virtual prototypes enable the model to accommodate future updates, serving as proxies dispersed throughout the embedding space to enhance the classifier’s performance during inference. SAVC [16] innovates by distinguishing new classes from existing ones through the incorporation of virtual classes into supervised contrastive learning. The virtual classes, generated through predefined transformations, serve not only as placeholders for unseen categories within the representation space but also enrich it with diverse semantic information. These virtual class methods demonstrate the importance of reserving embedding space for new classes and enhancing the separability of embeddings.
NC-FSCIL [12] sets up classifiers with an Equiangular Tight Frame (ETF) structure that considers future classes, and it induces class alignment with the corresponding classifier to mitigate the classifier misalignment issue during the incremental process. OrCo [25] maximizes the boundaries between classes and reserves space for subsequent incremental data by introducing perturbed prototypes in the feature space and maintaining orthogonality between classes. These methods of pre-allocated prototypes highlight the importance of prior space allocation and preventing the misalignment of old categories in new sessions.
FACT occupies embedding space through virtual prototypes, while SAVC generates negative samples for contrastive supervised learning via data augmentation to preserve embedding space. Additionally, contrastive learning enhances the model’s ability to generalize, which is crucial in FSL. NC-FSCIL assigns fixed prototypes to categories randomly before training and uses prototype alignment as the supervisory signal for the model. SAAN is a more flexible and semantically aligned learning strategy based on similarity estimation for spatial pre-allocation and dynamic adjustment of prototypes.
@@ 3.1 Problem Setting
In FSCIL, the model is fed with a sequence of training dataset $\left\{\mathcal{D}_{train}^{t}\right\}_{t=0}^{T}$, where $\mathcal{D}_{train}^{t}$ is the training set at the session $t$ and $T$ is the number of all sessions. $\mathcal{D}_{train}^{t}=\left\{(\mathbf{x}_{i},y_{i})\right\}$, where $\mathbf{x}_{i}$ is the $i$-th training sample and $y_{i}$ is the $i$-th label. $\left|\mathcal{D}_{train}^{t}\right|$ denotes the number of the training samples at session $t$. Each session $t$ has a distinct label space $\mathcal{Y}^{t}$, which means that $\mathcal{Y}^{t}\cap\mathcal{Y}^{t^{\prime}}=\varnothing$ for any $t\neq t^{\prime}$. $\left|\mathcal{Y}^{t}\right|$ denotes the number of the classes at session $t$. In the base session, $\left|\mathcal{D}_{train}^{0}\right|$ is equal to the number of all the samples whose labels $\in\mathcal{Y}^{0}$ in the dataset. For example, on CIFAR100 [50], there are 500 training samples for each class at the base session. However, at the incremental sessions, we only can access a limited number of training samples. The training data is organized in an $N$-way $K$-shot format which means there are $N$ classes, and each class contains $K$ training samples at each incremental session. For example, on CIFAR100, each incremental session includes 5 classes, with each class having 5 samples, meaning $N=5$ and $K=5$. At each session $t$, the model is tested on all test samples whose labels $\in(\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t})$.
@@ 3.2 The Model Framework
where ${n^{t}_{j}}$ is the number of samples of class $j$ and $\mathbb{I}\left(\cdot\right)$ is the indicator function. As indicated by Eq. 1, the NCM classifier, which classifies based on the nearest neighbor principle by measuring the cosine similarity between sample embeddings and representative points, completely disregards the norm information of the sample embeddings.
Algorithm 1 The training and classification process for the feature-freezing base approach.
1:The CNN $f(\cdot)$, The classifier $W$, The training set sequence $\left\{\mathcal{D}_{train}^{t}\right\}_{t=0}^{T}$, The model learning rate $\mu$
2:The updated CNN $f(\cdot)$, The updated classifier $W$
3:for each training dataset $\mathcal{D}_{train}^{t}$ do
4:     Expand the dimension of $W$ from $d\times|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t-1}|$ to $d\times|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|$
5:     if $t>0$ then
6:         Freeze most layers of the CNN for fine-tuning preparation
7:     end if
8:     while $\mathcal{L}_{ce}$ not converges do
9:         for each batch $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{m}\}\in\mathcal{D}_{train}^{t}$ do
10:              $e\leftarrow e+1$
11:              Compute $\mathcal{L}_{ce}$ on $\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{m}\}$
12:              $\theta_{f}^{e+1}\leftarrow\theta_{f}^{e}-\mu\frac{\partial\mathcal{L}_{ce}}{\partial\theta_{f}^{e}}$
13:              $W^{e+1}\leftarrow W^{e}-\mu\frac{\partial\mathcal{L}_{ce}}{\partial W^{e}}$
14:         end for
15:     end while
16:     Recalculate $W$ by Eq. 2 to replace fully connected layers by NCM
17:     Classify all the seen samples using the continuously expanding $W$ and the continuously updating $f$ through Eq. 1
18:end for
19:return $f(\cdot)$, $W$
@@ 3.3 Learning of the Baseline Model
The optimization goal of the feature extractor freezing paradigm baseline is to minimize $\mathcal{L}_{ce}(\phi(\mathbf{x}),y)$, where $\phi(\mathbf{x})=W^{\top}f(\mathbf{x})$ is the model composed of the feature extractor and the linear classifier, and $\mathcal{L}_{ce}(\cdot)$ is the cross-entropy loss function. Here, $\phi(\mathbf{x})\in\mathbb{R}^{|\mathcal{Y}|\times 1}$, $W\in\mathbb{R}^{d\times|\mathcal{Y}|}$, $f(\mathbf{x})\in\mathbb{R}^{d\times 1}$, and $d$ is the feature dimension. After the base session, most layers of the CNN are frozen and fine-tuned. The NCM classifier expands by $W=\{\boldsymbol{w}^{0}_{1},\boldsymbol{w}^{0}_{2},\cdots,\boldsymbol{w}^{t}_{|\mathcal{Y}^{t}|-1},\boldsymbol{w}^{t}_{|\mathcal{Y}^{t}|}\}$. $\boldsymbol{w}^{t}$ is continuously expanded during the incremental process based on new data. Although the feature-freezing base approach is simple, it outperforms many FSCIL methods, making it a strong baseline model. The training and classification algorithm depicting the feature-freezing base approach is shown in Alg. 1.
Figure 2: Illustration of CCSA. CCSA sets orthogonal class centers and assigns these centers to different categories to allocate the embedding space. CCSA uses the well-designed cosine center loss to guide embeddings into the corresponding subspace, achieving the effect of feature alignment. The centers are adjusted based on semantics in the base session, and newly inserted classes are allocated to the remaining space during the incremental sessions.
@@ 4.1 Class-Center Guided Embedding Space Allocation
The baseline model trained solely with cross-entropy loss focuses only on the current sample during training and utilizes the entire embedding space, rendering it challenging for the model to assimilate new class knowledge [18, 19]. The embedding space consists of all possible feature vectors. This space not only needs to accommodate the feature vectors of the current samples but also needs to reserve capacity for future classes. Inspired by the center loss [51], we mitigate this challenge by establishing learnable class centers (prototypes) and guiding sample embeddings toward their corresponding class centers. Unlike center loss, CCSA constrains the allocation of center positions and the update of centers to achieve the goal of constraining the model’s utilization of the embedding space and providing balanced space for new classes. The overall process of CCSA is shown in Fig. 2.
@@ 4.1.1 Space allocation
We first set $(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)$ orthogonal class centers and assign them one-to-one based on their distance from the class means. In the real world, considering the total number of classes is unknown and continuously growing, we just set $d$ orthogonal centers, which is the maximum number of orthogonal centers that can be generated in a $d$-dimensional feature space. To ensure that the total distance between all class centers and their corresponding class means is minimized, we use the Hungarian algorithm [52, 53] to assign the class centers to different categories. The Hungarian algorithm is used to solve the assignment problem in a bipartite graph, aiming to find an optimal matching that minimizes the total cost. Here, the cost is defined by cosine distance, because the fully connected layer in a neural network performs dot product operations, and some literature [54, 55] suggests that dot products can introduce bias due to different sample sizes. Cosine distance, on the other hand, can eliminate this bias by removing the influence of vector norms. We define an $(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)\times(|\mathcal{Y}^{0}\cup\mathcal{Y}^{1}\cup\dots\cup\mathcal{Y}^{t}|)$ cost matrix $D=[c_{i,j}]$, where $c_{i,j}$ represents the cosine distance from $i$-th class center to the $j$-th class. Our objective is to find a one-to-one matching that minimizes the total cost. Mathematically, this can be written as:
$\min_{\sigma}\sum_{i=1}^{n}c_{i,\sigma(i)},$